OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x - 2y = -4 2x - y = 1

OpenStudy (anonymous):

(2,3)

OpenStudy (anonymous):

x - 2y = -4 2x - y = 1. First step = times -2 to first equation then add first equation to 2nd to get rid of x. You get -x2 +4y = 8 2x -y = 1 now add them and your 2nd equation should become 3y = 9 now you can solve y, 3y = 9 y = 9/3 y = 3 now substitute y in either first or 2nd equation, 2x-(3) = 1 2x = 1 + 3 2x = 4 x = 4/2 x = 2 therefore answer for (x,y) = (2,3)